OpenStudy (anonymous):

What are the factors of f(x)? f(x) = x^3 + 5x^2 - x - 5

OpenStudy (anonymous):

x^3+5x^2-x-5 =x^2(x+5)-(x+5) =(x^2-1)(x+5) =(x+1)(x-1)(x+5) Does this help?

OpenStudy (dumbcow):

factor by grouping ^^